OpenStudy (voltron21):

HELP ME!!!!!!!!! Calculus! Do a delta-epsilon proof for:

OpenStudy (voltron21):

\[\lim_{x \rightarrow 4}\frac{ x^2-16 }{ x-4} =8\]

OpenStudy (anonymous):

Do you even need to do a delta epsilon proof? The top is (x-4)(x+4), so you're simply asked for the lim x->4 of x+4. That's pretty clearly 8...

OpenStudy (anonymous):

@elementalmagicks Wow you're confused, lol.

OpenStudy (voltron21):

i know but my teacher wants us to do it using a delta-epsilon proof :/

OpenStudy (anonymous):

\[ |x-4|<\delta\implies \left|\frac{x^2-16}{x-4}-8\right| <\epsilon \]

OpenStudy (anonymous):

I'm confused apparently. wio hath spoken.

OpenStudy (anonymous):

@voltron21 What is stopping you from finishing the proof?

OpenStudy (voltron21):

@wio im as confused with this whole concept as @elementalmagicks

OpenStudy (anonymous):

@elementalmagicks Yeah you are. There is a difference between a "proof" and just "evaluating"

OpenStudy (anonymous):

\[ |x-4|<\delta\implies \left|\frac{x^2-16}{x-4}-8\right| <\epsilon \]Becomes: \[ |x-4|<\delta\implies \left|\frac{(x-4)(x+4)}{x-4}-8\right| <\epsilon \]Since it is a limit, we assume \(x\neq 4\) so we have:\[ |x-4|<\delta\implies \left|(x+4)-8\right| <\epsilon \]

OpenStudy (anonymous):

@voltron21 Is it making sense yet?

OpenStudy (voltron21):

@wio kind of, so u set the restrictions and now you're simplifying

OpenStudy (anonymous):

I suppose I feel silly giving an "epsilon delta" proof when delta can just be chosen to be epsilon.

OpenStudy (anonymous):

Typically you have to choose more creatively, but this apparently was just asking you for the basic mechanics.

OpenStudy (anonymous):

The idea of epsilon delta proofs is that epsilon is fixed and then you pick a delta such that the implications that wio was been writing as satisfied.

OpenStudy (anonymous):

are*

OpenStudy (anonymous):

@voltron21 Yes, and with these sorts of proofs you generally want to get to something of the form: \[ |x-a| <\delta \implies g(x)|x-a|<\epsilon \] In this case \(g(x)=1\) so it is relatively easy \(\delta \sim \epsilon\). If \(g(x)=c\) (a constant) we'd pick \(\delta \sim \epsilon/ c\)

OpenStudy (voltron21):

@wio so is that it? is that the proof?

OpenStudy (anonymous):

I'm sort of explaining the methodology for the proof.

OpenStudy (voltron21):

@wio so ur not done?lol

OpenStudy (anonymous):

The proof was done a while ago.

OpenStudy (anonymous):

Once you get to:\[ |x-4|<\delta \implies |x-4|<\epsilon \]You just say \(\delta \sim \epsilon\) and you're done.

OpenStudy (voltron21):

@wio i thought i was supposed to get to the whole thing =epsilon

OpenStudy (anonymous):

You have to basically let \(\delta\) be a function of \(\epsilon\), which we did.

OpenStudy (voltron21):

@wio what does that little squiggle mean... sorry i dont understand math talk

OpenStudy (voltron21):

@wio and where did the 8 go?

OpenStudy (anonymous):

Don't worry about it, it doesn't have any special meaning.

OpenStudy (anonymous):

\(=\) here is fine as well

OpenStudy (voltron21):

@wio oh so it just means =

OpenStudy (anonymous):

\(\sim\) means in the ballpark of, technically delta doesn't have to equal epsilon since it can be less.

OpenStudy (voltron21):

@wio so it means about ok that makes sense but i still don't understand what happened to the 8

OpenStudy (anonymous):

\[ x+4-8=x-4 \]

OpenStudy (voltron21):

@wio oh lol got ya i just didnt see that...thanks!